The supplement exceeds an angle by 122 degrees, Find the supplement
Found 2 solutions by rfer, ewatrrr:
Answer by rfer(16322)

x+x+122=180
2x=58
x=29 degrees
x+122=151 degrees

Answer by ewatrrr(24785)

Hi,
Note: the sum of supplementary angles is 180
Let (x be the supplement aked for. (x - 122) the other angle.
.
x + (x - 122) = 180
.
2x - 122 = 180
.
2x = 302
x = 151 degrees
.
check you answer
151 + 29 = 180